OpenStudy (anonymous):

Let A = (0,0), B = (8,0), and C =(x,x). Find x, given that BC = 7

jimthompson5910 (jim_thompson5910):

Use the distance formula d = sqrt((x2-x1)^2+(y2-y1)^2) 7 = sqrt((x-8)^2+(x-0)^2) 7 = sqrt((x-8)^2+x^2) 7^2 = (x-8)^2+x^2 49 = (x-8)^2+x^2 49 = x^2-16x+64+x^2 49 = 2x^2-16x+64 0 = 2x^2-16x+64 - 49 2x^2-16x+15 = 0 I'll let you finish.

OpenStudy (anonymous):

Alright that's what I did originally but when I factored 2x^2-16x+15 = 0 and got x = 4 +- srqt(17/2) I thought I did something wrong. Thanks

jimthompson5910 (jim_thompson5910):

you should get \[\Large x = \frac{8\pm\sqrt{34}}{2}\]

OpenStudy (anonymous):

Ok I'll resolve this one